Electronics, according to Ohm's law, E = IR, where "E" is
voltage (measured in volts), "I" is current (measured in amps), and
"R" is resistance (measured in ohms).
What is the voltage across an 8-ohm resistor with a current of 1.5
amps flowing through it?

Answer by indra89811(24)

R=8 ohm
I=1.5 amp
V=I*R=1.5*8=12 volt
Ans: 12 volt
-----------------Peace--------------- :)